Find the zeroes of the function without using technology:
y=(3^-2(x-4)) -sqrt27

Answer by stanbon(75887)

Find the zeroes of the function without using technology:
y=(3^-2(x-4)) -sqrt27
-------------
Let y =0 and solve for "x"::
3^[-2(x-4)]= sqrt27
3^[-2(x-4)]= 3^(3/2)
-2(x-4) = (3/2)
x-4 = -3/4
x = 3 1/4 or 13/4
====================
Cheers,
Stan H.